# Quantifying Uncertainty: All We Need is the Bootstrap?

### v2

@@ Abstract
Quantifying uncertainty through standard errors, confidence intervals, hypothesis tests, and related measures is a fundamental aspect of statistical practice. However, these techniques involve a variety of methods, mathematical formulas, and underlying concepts, which can be complex. Could the non-parametric bootstrap, known for its simplicity and general applicability, serve as a universal alternative? In this study, we address this question through a review of existing literature and a simulation analysis of one- and two-sided confidence intervals across varying sample sizes, confidence levels, data-generating processes, and statistical functionals. Our findings indicate that the double bootstrap consistently performs best and is a promising alternative to traditional methods used for common statistical tasks. These results suggest that the bootstrap, particularly the double bootstrap, could simplify statistical education and practice without compromising effectiveness.
@@ 1 Introduction
In this paper, we focus on the quantification of uncertainty. Standard errors, confidence intervals, and hypothesis tests are integral components of statistical practice, yet they typically involve advanced concepts—such as test statistics and sampling distributions—and encompass a wide array of methods. Among these, one method stands out: the bootstrap. The bootstrap offers several advantages over traditional techniques for quantifying uncertainty. It is conceptually straightforward, reinforces the fundamental role of sampling in statistics, allows direct interaction with estimates and their distributions, and can be applied to a wide range of practical tasks without requiring the mastery of new concepts or complex mathematical formulas. These characteristics give the bootstrap significant pedagogical value [1], positioning it as a strong candidate for a one-size-fits-all approach to quantifying uncertainty, particularly for practitioners with limited statistical training.
Historically, the use of the bootstrap was constrained by computational limitations. With advancements in computing power, this is no longer a concern, yet introductory applied statistics textbooks and courses rarely place the bootstrap at the forefront. In our experience, this is largely due to the inertia of established practices, which are supported by more extensive instructional resources and software tools—resources that bootstrapping has historically lacked [1], and to some extent still does. Moreover, there is currently insufficient empirical evidence to convincingly establish the bootstrap as a comprehensive, one-size-fits-all solution for quantifying uncertainty.
@@ 1.1 Related work
Theoretical cases where bootstrap fails are pathological with limited implications for practice (see [2] and [3, Ch.2.6]). Large sample properties of common bootstrap approaches have also been established (see [2] for a summary). Unfortunately, large sample theory is not always a reliable predictor of finite sample performance, so we have to turn to empirical work.
Tables 1.1 and 1.1 summarize empirical studies on bootstrap methods. Early research primarily focused on the Pearson correlation and the sample mean, with subsequent work extending to other functionals, particularly quantiles and regression parameters. In this review, we concentrate on non-parametric bootstrap techniques and methods that are widely taught and applied in practice, referred to hereafter as baseline methods. These include, for example, Fisher confidence intervals (CIs) for the Pearson correlation and t-intervals for the sample mean.
@@ 2 Simulation Study
We ran an experiment for each possible combination of sample size $n$ from {4, 8, 16, 32, 64, 128, 256}, intervals $(-\infty,\alpha)$ with endpoints $\alpha$ from {0.025, 0.05, 0.25, 0.75, 0.95, 0.975}, statistical functional from {mean, median, standard deviation, $5^{th}$ and $95^{th}$ percentile, and Pearson correlation}, and data generating process from
@@
  • Laplace with $\mu=0$ and $b=1$, and
  • bi-normal with $\mu=\begin{bmatrix}1\\ 1\end{bmatrix}$ and $\Sigma=\begin{bmatrix}2&0.5\\ 0.5&1\end{bmatrix}$.
@@
Note that the bi-normal and Pearson correlation appear only in combination with each other.
@@
Our experimental design was informed by a combination of related simulation studies, theoretical considerations, and computational constraints. The selected range of sample sizes is consistent with those commonly used in similar studies. However, we include sample sizes of 4 and 8, which are rarely examined. Although such small samples are uncommon in most statistical practice, they may hold practical relevance in fields where data is limited (for example, gene expression studies). For sample sizes of 256 and above, we anticipate that all methods will perform well from a practical standpoint, though these larger sample sizes can also become computationally prohibitive.
@@
The selected range of endpoints includes the 5% and 95% confidence levels and enables the construction of symmetric two-sided confidence intervals at 95% and 90% levels by combining the 0.025 and 0.975, or 0.05 and 0.95 endpoints. These levels are the most commonly used in practice and in almost all simulation studies. However, our choice of endpoints also permits the exploration of other confidence levels. It is important to note, from an evaluation standpoint, that considering one-sided coverage is crucial, as relying solely on two-sided coverage can be misleading (see Section 2.2 for details).
@@
The selected statistical functionals include the most commonly used functionals (mean, median, standard deviation, and correlation) and two extreme percentiles, which are less commonly used in practice but are known to pose challenges in quantifying uncertainty. A notable omission is the inclusion of model coefficients, such as those from the family of generalized linear mixed-effects models. However, for many of these models, and for most more complex models, the bootstrap is not only a viable alternative but the only option for quantifying uncertainty.
We selected the data-generating processes to cover the entire spectrum of difficulty, which is primarily characterized by the skewness and kurtosis of the underlying distribution. A notable omission in our study is the consideration of categorical data, specifically the estimation of proportions in binary data. Addressing categorical data would require a different experimental setup, as our chosen data-generating processes are incompatible with proportion estimation, and the selected statistical functionals do not apply to categorical data. We delegate the analysis of categorical data to future work. However, note that in all but the smallest sample sizes and most extreme proportions, the mean and corresponding quantification of uncertainty will perform well.
@@
We used $n_{\text{rep}}=10000$ replications for every experiment, to limit coverage standard error to $0.005$ in the worst case, and $B=\{10,100,1000\}$ bootstrap replications. More bootstrap replications is better, so we do not consider $B$ a dimension of the experiment. We only report results for $B=1000$, but we would reach the same conclusions with $B=100$. For $B=10$ the performance of bootstrap methods is noticeably worse. Recommendations [3, 8, 38] and choices of $B$ in simulation studies also suggest that $B=1000$ is enough.
@@ 2.1 Methods
As a baseline for comparison, we include the CIs produced by the methods that are most commonly used in practice for that statistical functional. For the mean we use the t-based CIs from the commonly used t-test (t-test). For the median we use CIs from the Wilcoxon signed rank test (wilcoxon[39]. For standard deviation, we use the chi-squared CIs (chi-sq[40]. For Pearson correlation we use Fisher CIs (fisher) [41]. For quantiles we use parametric CIs based on normal assumption (q-par), non-parametric CI (q-nonpar) (see [7] for both), and the Maritz-Jarrett method (m-j) [42].
@@
Bootstrap sampling can be further categorized into parametric and non-parametric approaches. The parametric bootstrap assumes a specific distribution for the underlying population, $F$, and estimates the associated parameters from the observed data, $X=(x_{1},x_{2},\dots,x_{n})$. In contrast, the non-parametric bootstrap infers properties of $F$ by resampling directly from $X$ without imposing distributional assumptions. Given that there are $n^{n}$ possible samples from resampling, we typically restrict the process to $B$ independent bootstrap samples to maintain computational feasibility. This yields $B$ bootstrap samples $X^{*}_{1},X^{*}_{2},\dots,X^{*}_{B}$ and the bootstrap distribution of the parameter $\hat{\theta}^{*}=(\hat{\theta}^{*}_{1},\hat{\theta}^{*}_{2},\dots,\hat{\theta}^{*}_{B})$.
The assumptions underlying parametric bootstrap methods restrict their applicability or require additional user input, making them less suitable as a universal approach. Therefore, our experiments focus exclusively on non-parametric bootstrap methods. Below, we provide a brief overview of the bootstrap methods used for CI construction, along with references for further details. The source code for our implementation is also available (see Section 3).
@@ 2.1.1 Percentile Bootstrap (PB)
The percentile method is the original method proposed by Efron in [43] (for details, also see [38, chap. 13]). Even though multiple improvements have been made since, it is still one of the most popular bootstrap methods. The percentile CI for confidence level $\alpha$ is obtained by taking the $\alpha$-quantile of the bootstrap distribution:
@@
In all of our implementations of methods that use quantiles, we used the median-unbiased version of quantile calculation, recommended in [44].
@@ 2.1.2 Standard Bootstrap (B-n)
The standard method, sometimes also called the normal method, assumes that the bootstrap distribution is normal [38, chap. 13]:
@@ 2.1.3 Basic Bootstrap (BB)
In the basic bootstrap [38, chap. 13.4], sometimes called the reverse percentile method, we replace the observed bootstrap distribution, $\theta^{*}$ with $W^{*}=\theta^{*}-\hat{\theta}$. This results in
@@ 2.1.4 Smoothed Bootstrap (SB)
The smoothed bootstrap [3] gets its name from smoothing the bootstrap distribution. We implement smoothing with a normal kernel centered on 0. We determine kernel size with
@@ 2.1.7 Studentized Bootstrap (B-t)
But since the distribution of $T$ is not known, we need to approximate its percentiles from the bootstrap distribution. We do that by defining $T^{*}=(\hat{\theta}^{*}-\hat{\theta})/\hat{\sigma}^{*}$, where $\hat{\sigma}^{*}$ is obtained by doing another inner bootstrap sampling on each of the outer samples.
@@ 2.1.8 Double Bootstrap (DB)
The double bootstrap [45, chap. 3.11] corrects bias with another level of inner of bootstraps. We repeat the bootstrap procedure on each of the bootstrap samples to calculate the the percentage of times that the inner bootstrap functional is smaller than on the original sample. We want to take such a limit that $P\{\hat{\theta}\in(-\infty,\hat{\theta}_{double}[\alpha])\}=\alpha$, which is why we need to select the $\alpha$-th quantile of biases $\hat{b}^{*}$ for the adjusted level $\alpha_{\text{DB}}$. This leads to
@@ 2.2 Measuring the quality of confidence intervals
While we also focus on coverage, we measure and report results for interval length for two-sided CIs and the absolute distance from exact intervals for one-sided intervals. The exact interval for endpoint $\alpha$ and parameter $\theta$ is defined as $\hat{\theta}_{exact}[\alpha]=\hat{\theta}-\hat{\sigma}K^{-1}(1-\alpha)$, where $K$ is the cumulative distribution function of $\theta$ [46]. We approximate $\hat{\sigma}K^{-1}(1-\alpha)$ with 100000 samples.
Note that two-sided coverage can be misleading regarding a method’s coverage, because good two-sided coverage can be, and in practice often is, a result of substantial, but opposite errors in the two one-sided intervals (see [46] for an example). That is, while two-sided error can be studied from one-sided CIs, the converse is not true.
We argue that the practical meaning of coverage error depends on nominal coverage and is not symmetric. For example, 51% coverage at 50% nominal coverage is not the same as 96% coverage at 95% nominal coverage. And 85% coverage is not the same as 95% coverage at 90% nominal coverage. To aggregate results and for a threshold-based criterion that can be applied to all confidence levels, we propose a a novel criterion based on the Kullback-Leibler divergence. That is, we measure information loss if we assume nominal coverage $\pi$ when actual coverage is $p$:
@@
For a threshold-based criterion of what we consider good enough we modify Bradley’s criterion $|p-\pi|<\frac{\min(\pi,1-\pi)}{k}$, where $\pi$ is the nominal coverage [47]. Common choices for $k$ are 10 (stringent), 4 (intermediate), 2 (liberal), and 0.75 (very liberal). Note that the intermediate and very liberal were introduced by [48]. In related work on the bootstrap, only [37] uses a good enough criterion - that actual coverage should lie between 92.5% and 97.5% when nominal coverage is 95% (this is based on the work of [47]).
At $95\%$ nominal coverage, the KL divergences for the Bradley lower bounds $94.5\%$ ($k=10$), $93.8\%$ ($k=4$), $92.5\%$ ($k=2$), and $88.3\%$ ($k=0.75$) are approximately $0.0004$, $0.0020$, $0.0083$, and $0.0503$. We adopt $\text{KL}(0.945,0.95)$ as the stringent criterion and a factor of 5 as an order of magnitude worse/better performance. For nominal coverage 95% this leads to criteria very similar to Bradley’s: $(93.8,96.2)$ vs $(93.9,96.1)$ (intermediate), $(92.5,97.5)$ vs $(92.4,97.3)$ (liberal), and $(88.3,101.7)$ vs $(88.6,99.4)$ (very liberal). However, for nominal coverage further away from 95% our approach gives produces sensible criteria and does not produce endpoints outside of the unit interval.
@@ 3 Results
The simulation study consists of 1302 combinations and it is infeasible to list them all. We will focus on identifying if a bootstrap method is a viable one-size-fits-all approach. Inevitably, some details that are relevant to the reader might be left out. We prepared a visualization tool to browse all of the the results of the experiments zrimseku.github.io/bootstrap-simulation/ (see Figure 2).
The raw results (aggregated over 10000 replications) and the source code for the tool, simulation framework, pre-processing, and analysis can be found here: github.com/zrimseku/bootstrap-simulation. The source code can be used to generate the full non-aggregated results. The library with all the bootstrap methods can be found here: github.com/zrimseku/bootstrap-ci.

### v3

@@ Abstract
{A critical literature review and comprehensive simulation study is used to show that (a) non-parametric bootstrap is a viable alternative to commonly taught and used methods in basic estimation tasks (mean, variance, quartiles, correlation) and (b), contrary to recommendations in most related work, double bootstrap performs better than BCa.} Quantifying uncertainty through standard errors, confidence intervals, hypothesis tests, and related measures is a fundamental aspect of statistical practice. However, these techniques involve a variety of methods, mathematical formulas, and underlying concepts, which can be complex. Could the non-parametric bootstrap, known for its simplicity and general applicability, serve as a universal alternative? This paper addresses this question through a review of the existing literature and a simulation analysis of one- and two-sided confidence intervals across varying sample sizes, confidence levels, data-generating processes, and statistical functionals. Results show that the double bootstrap consistently performs best and is a promising alternative to traditional methods used for common statistical tasks. These results suggest that the bootstrap, particularly the double bootstrap, could simplify statistical education and practice without compromising effectiveness.
@@ 1 Introduction
This paper focuses on the quantification of uncertainty. Standard errors, confidence intervals, and hypothesis tests are integral components of statistical practice, yet they typically involve advanced concepts—such as test statistics and sampling distributions—and encompass a wide array of methods. Among these, one method stands out: the bootstrap. The bootstrap offers several advantages over traditional techniques for quantifying uncertainty. It is conceptually straightforward, reinforces the fundamental role of sampling in statistics, allows direct interaction with estimates and their distributions, and can be applied to a wide range of practical tasks without requiring the mastery of new concepts or complex mathematical formulas. These characteristics give the bootstrap significant pedagogical value [1], positioning it as a strong candidate for a one-size-fits-all approach to quantifying uncertainty, particularly for practitioners with limited statistical training.
Historically, the use of the bootstrap was constrained by computational limitations. With advancements in computing power, this is no longer a concern, yet introductory applied statistics textbooks and courses rarely place the bootstrap at the forefront. This is largely due to the inertia of established practices, which are supported by more extensive instructional resources and software tools—resources that bootstrapping has historically lacked [1], and to some extent still does. Moreover, there is currently insufficient empirical evidence to convincingly establish the bootstrap as a comprehensive, one-size-fits-all solution for quantifying uncertainty.
@@ 1.1 Related work
Theoretical cases where bootstrap fails are pathological with limited implications for practice (see [2] and [3, Ch.2.6]). Large sample properties of common bootstrap approaches have also been established (see [2] for a summary). Unfortunately, large sample theory is not always a reliable predictor of finite sample performance, so empirical work is required.
Tables 1.1 and 1.1 summarize empirical studies on bootstrap methods. Early research primarily focused on the Pearson correlation and the sample mean, with subsequent work extending to other functionals, particularly quantiles and regression parameters. In this review, the focus is on non-parametric bootstrap techniques and methods that are widely taught and applied in practice, referred to hereafter as baseline methods. These include, for example, Fisher confidence intervals (CIs) for the Pearson correlation and t-intervals for the sample mean.
@@ 2 Simulation Study
The experiment runs across each possible combination of sample size $n$ from {4, 8, 16, 32, 64, 128, 256}, intervals $(-\infty,\alpha)$ with endpoints $\alpha$ from {0.025, 0.05, 0.25, 0.75, 0.95, 0.975}, statistical functional from {mean, median, standard deviation, $5^{th}$ and $95^{th}$ percentile, and Pearson correlation}, and data generating process from
@@
  • Laplace with $\mu=0$ and $b=1$,
  • Bernoulli with $p=0.5$,
  • Bernoulli with $p=0.9$, and
  • bivariate normal with $\mu=\begin{bmatrix}1\\ 1\end{bmatrix}$ and $\Sigma=\begin{bmatrix}2&0.5\\ 0.5&1\end{bmatrix}$.
@@
Note that the bivariate normal and Pearson correlation appear only in combination with each other and the two Bernoulli distributions appear only in combination with the mean.
@@
The experimental design was informed by a combination of related simulation studies, theoretical considerations, and computational constraints. The selected range of sample sizes is consistent with those commonly used in similar studies. However, sample sizes of 4 and 8, which are rarely examined, are included. Although such small samples are uncommon in most statistical practice, they may hold practical relevance in fields where data is limited (for example, gene expression studies). For sample sizes of 256 and above, all methods should perform well from a practical standpoint, though these larger sample sizes can also become computationally prohibitive.
@@
The selected range of endpoints includes the 5% and 95% confidence levels and enables the construction of symmetric two-sided confidence intervals at 95% and 90% levels by combining the 0.025 and 0.975, or 0.05 and 0.95 endpoints. These levels are the most commonly used in practice and almost all simulation studies. However, this choice of endpoints also permits the exploration of other confidence levels. It is important to note, from an evaluation standpoint, that considering one-sided coverage is crucial, as relying solely on two-sided coverage can be misleading (see Section 2.2 for details).
@@
The selected statistical functionals include the most commonly used functionals (mean, median, standard deviation, and correlation) and two extreme percentiles, which are less widely used in practice but are known to pose challenges in quantifying uncertainty. A notable omission is the inclusion of model coefficients, such as those from the family of generalized linear mixed-effects models. However, for many of these models, and most more complex models, the bootstrap is not only a viable alternative but the only option for quantifying uncertainty.
@@
Every experiment is replicated $n_{\text{rep}}=10000$ times, to limit coverage standard error to $0.005$ in the worst case, and $B=\{10,100,1000\}$ bootstrap replications. More bootstrap replications is better, so $B$ is not considered as a dimension of the experiment. Results are only reported for $B=1000$, but $B=100$ would result in the same main conclusions. For $B=10$ the performance of bootstrap methods is noticeably worse. Recommendations [3, 8, 38] and choices of $B$ in simulation studies also suggest that $B=1000$ is sufficient.
@@ 2.1 Methods
CIs produced by the methods that are most commonly used in practice for that statistical functional are included as a baseline for comparison. For the mean, the t-based CIs from the commonly used t-test (t-test). For the mean of the two Bernoulli distributions, the Clopper-Pearson (c-p[39] and Agresti-Coull (a-c[40] intervals. For the median, CIs from the Wilcoxon signed rank test (wilcoxon[41]. For standard deviation, chi-squared CIs (chi-sq[42]. For Pearson correlation Fisher CIs (fisher) [43]. For quantiles, parametric CIs based on normal assumption (q-par), non-parametric CI (q-nonpar) (see [7] for both), and the Maritz-Jarrett method (m-j) [44].
@@
Bootstrap sampling can be further categorized into parametric and non-parametric approaches. The parametric bootstrap assumes a specific distribution for the underlying population, $F$, and estimates the associated parameters from the observed data, $X=(x_{1},x_{2},\dots,x_{n})$. In contrast, the non-parametric bootstrap infers properties of $F$ by resampling directly from $X$ without imposing distributional assumptions. Given that there are $n^{n}$ possible samples from resampling, the process is typically restricted to $B$ independent bootstrap samples to maintain computational feasibility. This yields $B$ bootstrap samples $X^{*}_{1},X^{*}_{2},\dots,X^{*}_{B}$ and the bootstrap distribution of the parameter $\hat{\theta}^{*}=(\hat{\theta}^{*}_{1},\hat{\theta}^{*}_{2},\dots,\hat{\theta}^{*}_{B})$.
The assumptions underlying parametric bootstrap methods restrict their applicability or require additional user input, making them less suitable as a universal approach. Therefore, the experiments focus exclusively on non-parametric bootstrap methods. Below is a brief overview of the bootstrap methods used for CI construction, along with references for further details. The source code is also available (see Section 3).
@@ 2.1.1 Percentile Bootstrap (PB)
The percentile method is the original method proposed by Efron in [45] (for details, also see [38, chap. 13]). Multiple improvements have been made since, but percentile remains one of the most popular bootstrap methods. The percentile CI for confidence level $\alpha$ is obtained by taking the $\alpha$-quantile of the bootstrap distribution:
@@
All implementations of methods that use quantiles, use the median-unbiased version of quantile calculation, recommended in [46].
@@ 2.1.2 Standard Bootstrap (B-n)
The standard method (sometimes the normal method), assumes that the bootstrap distribution is normal [38, chap. 13]:
@@ 2.1.3 Basic Bootstrap (BB)
In the basic bootstrap [38, chap. 13.4], sometimes called the reverse percentile method, the observed bootstrap distribution $\theta^{*}$ is replaced with $W^{*}=\theta^{*}-\hat{\theta}$. This results in
@@ 2.1.4 Smoothed Bootstrap (SB)
The smoothed bootstrap [3] gets its name from smoothing the bootstrap distribution. Smoothing is implemented with a normal kernel centered on 0 and kernel size is determined with
@@ 2.1.7 Studentized Bootstrap (B-t)
But since the distribution of $T$ is not known, its percentiles have to be approximated from the bootstrap distribution. That is done by defining $T^{*}=(\hat{\theta}^{*}-\hat{\theta})/\hat{\sigma}^{*}$, where $\hat{\sigma}^{*}$ is obtained by doing another inner bootstrap sampling on each of the outer samples.
@@ 2.1.8 Double Bootstrap (DB)
The double bootstrap [47, chap. 3.11] corrects bias with another level of inner of bootstraps. The bootstrap procedure is repeated on each of the bootstrap samples to calculate the the percentage of times that the inner bootstrap functional is smaller than on the original sample. A limit is required such that $P\{\hat{\theta}\in(-\infty,\hat{\theta}_{double}[\alpha])\}=\alpha$, which is why the $\alpha$-th quantile of biases $\hat{b}^{*}$ is selected for the adjusted level $\alpha_{\text{DB}}$. This leads to
@@
Note that percentile bootstrap is used on the inner and outer bootstrap, but the double (iterated) bootstrap allows for any bootstrap method.
@@ 2.2 Measuring the quality of confidence intervals
This study also focuses on coverage, but also measures and reports results for interval length for two-sided CIs and the absolute distance from exact intervals for one-sided intervals. The exact interval for endpoint $\alpha$ and parameter $\theta$ is defined as $\hat{\theta}_{exact}[\alpha]=\hat{\theta}-\hat{\sigma}K^{-1}(1-\alpha)$, where $K$ is the cumulative distribution function of $\theta$ [48]. Note that $\hat{\sigma}K^{-1}(1-\alpha)$ is approximated with 100000 samples.
Note that two-sided coverage can be misleading regarding a method’s coverage because good two-sided coverage can be, and in practice often is, a result of substantial, but opposite errors in the two one-sided intervals (see [48] for an example). That is, while two-sided error can be studied from one-sided CIs, the converse is not true.
The practical meaning of coverage error depends on nominal coverage and is not symmetric. For example, 51% coverage at 50% nominal coverage is not the same as 96% coverage at 95% nominal coverage. And 85% coverage is not the same as 95% coverage at 90% nominal coverage. To aggregate results and for a threshold-based criterion that can be applied to all confidence levels, a novel criterion is proposed, based on the Kullback-Leibler divergence. That is, it is based on measuring information loss if nominal coverage $\pi$ is assumed when actual coverage is $p$:
@@
For a threshold-based criterion of what is considered good enough Bradley’s criterion $|p-\pi|<\frac{\min(\pi,1-\pi)}{k}$ is modified, where $\pi$ is the nominal coverage [49]. Common choices for $k$ are 10 (stringent), 4 (intermediate), 2 (liberal), and 0.75 (very liberal). Note that the intermediate and very liberal were introduced by [50]. In related work on the bootstrap, only [37] uses a good enough criterion - that actual coverage should lie between 92.5% and 97.5% when nominal coverage is 95% (this is based on the work of [49]).
At $95\%$ nominal coverage, the KL divergences for the Bradley lower bounds $94.5\%$ ($k=10$), $93.8\%$ ($k=4$), $92.5\%$ ($k=2$), and $88.3\%$ ($k=0.75$) are approximately $0.0004$, $0.0020$, $0.0083$, and $0.0503$. The value $\text{KL}(0.945,0.95)$ is adopted as the stringent criterion and a factor of 5 as an order of magnitude worse/better performance. For nominal coverage 95% this leads to criteria very similar to Bradley’s: $(93.8,96.2)$ vs $(93.9,96.1)$ (intermediate), $(92.5,97.5)$ vs $(92.4,97.3)$ (liberal), and $(88.3,101.7)$ vs $(88.6,99.4)$ (very liberal). However, for nominal coverage further away from 95%, the proposed approach gives produces sensible criteria and does not produce endpoints outside of the unit interval.
@@ 3 Results
The simulation study consists of 1386 combinations of sample size, endpoint, and functional, and it is infeasible to list all of the results. The results first focus on identifying if a bootstrap method is a viable one-size-fits-all approach. Inevitably, some details that are relevant to the reader might be left out. A visualization tool is available to browse all of the results of the experiments zrimseku.github.io/bootstrap-simulation/ (see Figure 2).
The raw results (aggregated over 10000 replications) and the source code for the tool, simulation framework, pre-processing, and analysis can be found here: github.com/zrimseku/bootstrap-simulation. The source code can be used to generate the full non-aggregated results. The library with all the methods can be found here: github.com/zrimseku/bootstrap-ci.